Question 40 Use implicit differentiation to find dy/dx. 2xy = y 2 1 0 0 yax
Answer #1

2xy - y2 = 1 Differentiating with respect x, we get 2x +2y(1) - 2y dy : 0 x+ywy (x - y) + y = 0 (x-y) =-Y = dx (x- y) . = Y-X